OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (-4,2)and y Intercept (0,-30) Wright answer if this form (x1,y1)(x2,y2).

OpenStudy (anonymous):

i would also like to know how you did it

Nnesha (nnesha):

what's the vertex form of the parabola ?

OpenStudy (anonymous):

(y-k)=a(x-h)^2

OpenStudy (anonymous):

Bruh... okay here: http://openstudy.com/study#/updates/4fe123c7e4b06e92b8705343

Nnesha (nnesha):

hmm add k both sides so it would be \[\huge\rm y=a(x-h)^2+k\] vertex form where (h,k) is the vertex so substitute (h,k) for ( -4,2)

OpenStudy (anonymous):

so like (y-2)=a(x--4)^2

Nnesha (nnesha):

ye that would work so y-intercept is (0,-30) where x =0 and y =-30 so plugin \[\huge\rm -30-2=a(0-(-4))^2\] now solve for a

OpenStudy (anonymous):

so a =-2

OpenStudy (anonymous):

i appreciate your help btw

Nnesha (nnesha):

yes right now u can find x-intercept \[\huge \rm -30-2=-2(x-(-4))^2\] solve for x let me know if you don't know how to work it out

OpenStudy (anonymous):

x= 0,-8 Correct?

Nnesha (nnesha):

hmm plz show a little work...

OpenStudy (anonymous):

im using an online calculator because the only calculator we have broke im not quite shure how to do the steps myself without it.

Nnesha (nnesha):

hmm well do you want to know how to work it out without calculator ?

OpenStudy (anonymous):

couldent hurt mind if i get my notepad and pencil

OpenStudy (anonymous):

ok i got it

Nnesha (nnesha):

alright first we can multiply negative by (-4) -1 times -4 = 4 ( negative times negative = positive ) \[\huge \rm -30-2=-2(x+4))^2\] -30-2 = -32 \[\huge \rm -32=-2(x+4)^2\] now we should foil (x+4)^2 `(x+4)^2 `is same as (x+4)(x+4)

Nnesha (nnesha):

 now we should distribute 2nd parentheses by first and 2nd term of the 2nd parentheses (foil method ) so multiply left first term by top